Give an example of an event that has a probability of one?
If I drop an object, it will fall toward earth. Entropy. My pet betta, Gil Fynn, will die. Death and taxes
the answer ot this question is 3/10
Ummmm no.. that's not correct at all.
Each machine is an independent event. so the correct answer is 0.001..
.1 * .1 * .1 = 0.001
or 0.1%
What is the probability that a huge asteroid will hit us?
I just watched a video about this yesterday......by the end of the century the odds that an asteroid wipes out millions is 1/20,000. They are trying to find ways to prevent a collision, however.
Dress inapropriately and talk to strangers. It might lead to sex but it would still b napping if there was any force.
i know im rh neg and my 2 brothers are rh neg. So i guess we ar eok. i have 3 children myself, but i dont know their blood type.
you have a 1/15 chance to win. or you could say a 6.66 % chance.
What are the odds of being run over tomorrow?
50 50 chances depends if your crossing a road or walking causally and see a nice lady then bamm a car hits you
Acutally the chances of having a boy are 51% and a girl 49%. The odds stay pretty close to 50% for each child and only vary slightly. If you have had 2 or 3 boys, you are only about 2% to 6% more likely to have another boy. If you have had girls, you are slightly more likely to have a boy next.
What does theoretical probability mean?
It's the number of favorable outcomes & the number of possible outcomes.
What is the probability of getting a tail if a coin tossed once?
Since the coin only has two sides, the probability of getting either heads or tails in any one toss is 1 in 2, or 50%. 50 precent chance. Everytime you toss it, 50 percent.
If an event is certain to happen which number represents its probability?
The probability of an event that is certain to happen is 1.
You have a 4/52 chance of your card being an ace the first time and a 4/52 chance of your card being an ace the second time.
4/52 * 4/52 or (Simplified) 1/13 * 1/13 = 1/169
26/52 cards are black and 13/52 cards are diamonds so the total probability of either a diamond or a black card is 39/52 or 3/4
p = 0.75
it is something that is almost the same if you think
Once you have picked the first ball, you will have five left to pick from, one of which will be the same color as the first pick, and four of colors that you don't want. The probability of picking the second ball of the same color is then 1/5 or 0.20. P.S. It really isn't fair to have someone else do your homework for you.
How does probability differ from actuality?
Probability is the chance of some outcome while actuality is the realistic chance and actual outcome of an event.
The best way to think about this question might be to think about a slightly different one: what is the probability that a randomly-chosen item is neither rusted nor a bolt? Well in this case it would have to be an un-rusted nut, and there are (1/2)*150=75 of these. So the probability of choosing an un-rusted nut is equal to the number of un-rusted nuts over the total number of items in the drawer, or 75/200=3/8=0.375. Now, everything in the drawer is either an un-rusted nut or it isn't, so the probability that the item we choose belongs to one or the other of those sets is 1 (it happens with certainty). Thus the probability that we choose something other than an un-rusted nut (something that is rusted or a bolt) is the difference in the two probabilities, or 1-3/8=5/8=0.625.
To get the same answer directly, you'd have to notice that there 25 rusty bolts and 75 rusty nuts, plus 25 bolts that aren't rusty, and so you're interested in the probability of choosing one of these 125 items out of the full drawer of 200 items. This is 125/200=5/8=0.625 as before. You can't just add the number of rusty items (100) to the number of bolts (50) and divide this by 200, because you then double-count the rusty bolts (each one belongs to the rusty group and to the bolt group). A backwards way of answering the question would be to add the number of rusty items to the number of bolts, and then subtract the number of rusty bolts, and finally divide by 200. Subtracting out the number of rusty bolts makes sure that these aren't double counted, and so the answer is again 125/200=5/8=0.625. This last approach is related to a more general idea called the Inclusion-Exclusion Principle.
There are a total of 25 Marbles The chances are 3 out of 25 drawing a Red marble. 3/25 = 12% chance of drawing a red marble
If a die is rolled 60 times what is the probability of having a 2 appear at least 9 times?
It is 1 - 0.3120 = 0.6880, approx.
The probability that fewer than 8 out of 10?
To get more than 8 one needs to get 9 or 10 so there are 2/10 chances
The probability to get less than 8 is the opposite probability to the one before
so it's 1-2/10 = 8/10 = 4/5 = 0.8